# Scene-wise Adaptive Network for Dynamic Cold-start Scenes Optimization in CTR Prediction

### v1

@@ Abstract.
In the realm of modern mobile E-commerce, providing users with nearby commercial service recommendations through location-based online services has become increasingly vital. While machine learning approaches have shown promise in multi-scene recommendation, existing methodologies often struggle to address cold-start problems in unprecedented scenes: the increasing diversity of commercial choices, along with the short online lifespan of scenes, give rise to the complexity of effective recommendations in online and dynamic scenes. In this work, we propose Scene-wise Adaptive Network (SwAN), a novel approach that emphasizes high-performance cold-start online recommendations for new scenes. Our approach introduces several crucial capabilities, including scene similarity learning, user-specific scene transition cognition, scene-specific information construction for the new scene, and enhancing the diverged logical information between scenes. We demonstrate SwAN’s potential to optimize dynamic multi-scene recommendation problems by effectively online handling cold-start recommendations for any newly arrived scenes. More encouragingly, SwAN has been successfully deployed in Meituan’s online catering recommendation service, which serves millions of customers per day, and SwAN has achieved a 5.64% CTR index improvement relative to the baselines and a 5.19% increase in daily order volume proportion.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italyisbn: 978-1-4503-XXXX-X/18/06submissionid: 163ccs: Computing methodologies Probabilistic reasoning
@@ 3.3. Cross-scene Feature Representation (CFR)
where $E_{o}$ means the embedding of other features, and $+$ means element-wise addition ( A in Fig. 2).
@@ 4.3. Hyperparameters Experiments
Table 6. Model performance under different $\alpha$ ($\beta=0.001,\gamma=0.001$).
@@
Table 7. Model performance under different $\beta$ ($\alpha=1,\gamma=0.001$).
@@
Table 8. Model performance under different $\gamma$ ($\alpha=1,\beta=0.001$).
@@
Table 10. Model performance under different temperature coefficient $\tau$.
@@ 4.6. Model Complexity
We compared each model’s number of parameters, training time, and inference time. All experiments were conducted on NVIDIA Tesla A100 GPU, 80G RAM, and Intel(R) Xeon(R) Gold 5218 CPU servers. Table 14 shows that SwAN maintains a reasonable number of parameters and prediction time, indicating that it can be deployed online without incurring additional costs. Generally, we retrieve online data in real-time for training and update the model approximately every half hour.

### v3

@@ Abstract.
In the realm of modern mobile E-commerce, providing users with nearby commercial service recommendations through location-based online services has become increasingly vital. While machine learning approaches have shown promise in multi-scene recommendation, existing methodologies often struggle to address cold-start problems in unprecedented scenes: the increasing diversity of commercial choices, along with the short online lifespan of scenes, give rise to the complexity of effective recommendations in online and dynamic scenes. In this work, we propose Scene-wise Adaptive Network (SwAN  [^1] https://github.com/ChrisLiiiii/SwAN ), a novel approach that emphasizes high-performance cold-start online recommendations for new scenes. Our approach introduces several crucial capabilities, including scene similarity learning, user-specific scene transition cognition, scene-specific information construction for the new scene, and enhancing the diverged logical information between scenes. We demonstrate SwAN’s potential to optimize dynamic multi-scene recommendation problems by effectively online handling cold-start recommendations for any newly arrived scenes. More encouragingly, SwAN has been successfully deployed in Meituan’s online catering recommendation service, which serves millions of customers per day, and SwAN has achieved a 5.64% CTR index improvement relative to the baselines and a 5.19% increase in daily order volume proportion.
@@
journalyear: 2024copyright: acmlicensedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688115isbn: 979-8-4007-0505-2/24/10submissionid: 163ccs: Computing methodologies Probabilistic reasoning
@@ 3.3. Cross-scene Feature Representation (CFR)
where $E_{o}$ means the embedding of other features, and $+$ means element-wise addition.
@@ 4.3. Hyperparameters Experiments
Table 6. Model under different $\alpha$ ($\beta=0.001,\gamma=0.001$).
@@
Table 7. Model under different $\beta$ ($\alpha=1,\gamma=0.001$).
@@
Table 8. Model under different $\gamma$ ($\alpha=1,\beta=0.001$).
@@
Table 10. Model under different temperature coefficient $\tau$.
@@ 4.6. Model Complexity
All experiments were conducted on NVIDIA Tesla A100 GPU, 80G RAM, and Intel(R) Xeon(R) Gold 5218 CPU servers. Table 14 shows that SwAN maintains a reasonable number of parameters and prediction time. Generally, we retrieve online data in real-time for training and update the model approximately every half hour.
@@ 5. Conclusion
Acknowledgements.
This work was supported in part by the National Key Research and Development Program of China under Grant 2023YFB3307503, in part by the National Natural Science Foundation of China under Grant 62202025, in part by the Young Elite Scientist Sponsorship Program by CAST under Grant YESS20230566, in part by the CCF-Huawei Populus Grove Fund, in part by the Frontier Cross Fund Project of Beihang University, and in part by the Fundamental Research Fund Project of Beihang University.